# The Non Parametric Reconstruction of Binary Black Hole Mass Evolution from GWTC-4.0 Gravitational Wave Catalog

### v1

@@ 5 Conclusions
Looking forward, the dramatic increase in detection rates expected from next generation ground based detectors such as Cosmic Explorer (Reitze et al., 2019) and the Einstein Telescope (Branchesi et al., 2023) will extend BBH catalogs to higher redshift, enabling reconstruction of higher order evolutionary features with high statistical precision. Our non parametric methodology is well positioned to exploit these datasets, offering both robustness and flexibility in probing the cosmic history of black hole formation. As GW astronomy transitions into the precision era, such model independent frameworks will be essential for extracting the full astrophysical information encoded in BBH populations, ultimately linking GW observations to stellar evolution, galactic chemical enrichment, and the early universe.
@@
The gravitational wave catalogs used in this study are publicly available on Zenodo as part of the LVK data releases: GWTC-2.1 (link), GWTC-3 (link) and GWTC-4.0 (link).
Acknowledgments
The authors express their gratitude to Souradeep Pal for reviewing the manuscript and providing useful comments as a part of the LIGO publication policy. This work is part of the ⟨data|theory⟩ Universe Lab, supported by TIFR and the Department of Atomic Energy, Government of India. The authors express gratitude to the system administrator of the computer cluster of ⟨data|theory⟩ Universe Lab. This research is supported by the Prime Minister Early Career Research Award, Anusandhan National Research Foundation, Government of India. We thank the LIGO Scientific Collaboration, the Virgo Collaboration, and the KAGRA Collaboration for providing the gravitational-wave data through the Gravitational Wave Open Science Center (GWOSC). LIGO, funded by the U.S. National Science Foundation (NSF), and Virgo, supported by the French CNRS, Italian INFN, and Dutch Nikhef, along with contributions from Polish and Hungarian institutes. This collaborative effort is backed by the NSF’s LIGO Laboratory, a major facility fully funded by the National Science Foundation. The research leverages data and software from the Gravitational Wave Open Science Center, a service provided by LIGO Laboratory, the LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA. Advanced LIGO’s construction and operation receive support from STFC of the UK, Max Planck Society (MPS), and the State of Niedersachsen/Germany, with additional backing from the Australian Research Council. Virgo, affiliated with the European Gravitational Observatory (EGO), secures funding through contributions from various European institutions. Meanwhile, KAGRA’s construction and operation are funded by MEXT, JSPS, NRF, MSIT, AS, and MoST. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge the use of the following packages in this work: Numpy (Van Der Walt et al., 2011), Scipy (Jones et al., 2001–), Matplotlib (Hunter, 2007), Astropy (Robitaille et al., 2013) and emcee (Foreman-Mackey et al., 2013).

### v2

@@ V Conclusions
Looking forward, the increase in detection rates expected from next generation ground based detectors such as Cosmic Explorer (Reitze et al., 2019) and the Einstein Telescope (Branchesi et al., 2023) will extend BBH catalogs to higher redshift, enabling reconstruction of higher order evolutionary features with high statistical precision. Our non parametric methodology is well positioned to exploit these datasets, offering both robustness and flexibility in probing the cosmic history of black hole formation. As GW astronomy transitions into the precision era, such model independent frameworks will be essential for extracting the full astrophysical information encoded in BBH populations, ultimately linking GW observations to stellar evolution, galactic chemical enrichment, and the early universe.
Acknowledgments
The authors express their gratitude to Souradeep Pal for reviewing the manuscript and providing useful comments as a part of the LIGO publication policy. This work is part of the ⟨data|theory⟩ Universe Lab, supported by TIFR and the Department of Atomic Energy, Government of India. The authors express gratitude to the system administrator of the computer cluster of ⟨data|theory⟩ Universe Lab. This research is supported by the Prime Minister Early Career Research Award, Anusandhan National Research Foundation, Government of India. We thank the LIGO Scientific Collaboration, the Virgo Collaboration, and the KAGRA Collaboration for providing the gravitational-wave data through the Gravitational Wave Open Science Center (GWOSC). LIGO, funded by the U.S. National Science Foundation (NSF), and Virgo, supported by the French CNRS, Italian INFN, and Dutch Nikhef, along with contributions from Polish and Hungarian institutes. This collaborative effort is backed by the NSF’s LIGO Laboratory, a major facility fully funded by the National Science Foundation. The research leverages data and software from the Gravitational Wave Open Science Center, a service provided by LIGO Laboratory, the LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA. Advanced LIGO’s construction and operation receive support from STFC of the UK, Max Planck Society (MPS), and the State of Niedersachsen/Germany, with additional backing from the Australian Research Council. Virgo, affiliated with the European Gravitational Observatory (EGO), secures funding through contributions from various European institutions. Meanwhile, KAGRA’s construction and operation are funded by MEXT, JSPS, NRF, MSIT, AS, and MoST. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We acknowledge the use of the following packages in this work: Numpy (Van Der Walt et al., 2011), Scipy (Jones et al., 2001–), Matplotlib (Hunter, 2007), Astropy (Robitaille et al., 2013), pygtc (Bocquet & Carter, 2016), and emcee (Foreman-Mackey et al., 2013).
Appendix A Posterior Distributions of Evolution Coefficients
In this appendix we present the full posterior distributions of the non-parametric evolution coefficients inferred in our hierarchical Bayesian analysis. The intrinsic mass-redshift distribution is expanded as a Taylor series in redshift,
$\rm{p(m,z)=p_{0}(m)+p_{1}(m)(z-z_{\rm ref})+p_{2}(m)(z-z_{\rm ref})^{2}},$ (A1)
where $\rm{p_{0}(m)}$ represents the baseline mass distribution, while $\rm{p_{1}(m)}$ and $\rm{p_{2}(m)}$ capture the linear and quadratic redshift evolution of the population, respectively. We show the full posterior distributions and parameter correlations using corner plots. These visualizations provide a comprehensive view of parameter uncertainties, degeneracies, and the stability of the hierarchical inference across different binning choices.
Figures 6 and 7 show the posterior distributions for the primary mass coefficients using coarse ($\rm{\Delta M=10M_{\odot}}$) and fine ($\rm{\Delta M=5M_{\odot}}$) mass binning schemes, respectively. Figures 8 and 9 show the corresponding results for the secondary mass coefficients. The diagonal panels show the marginalized one-dimensional posterior distributions for each parameter, while the off-diagonal panels display the two-dimensional joint posteriors with $1\sigma$ and $2\sigma$ credible regions.
Figure 6: Corner plot showing the posterior distributions of the evolution coefficients for the primary black hole mass ($m_{1}$) using the coarse binning scheme ($\rm{\Delta M=10M_{\odot}}$). The parameters correspond to the Taylor expansion coefficients $\{p_{0}(m_{k}),p_{1}(m_{k}),p_{2}(m_{k})\}$. The diagonal panels show the marginalized one-dimensional posterior distributions, while the off-diagonal panels show the joint posterior distributions with $1\sigma$ and $2\sigma$ credible regions. This configuration contains 21 parameters corresponding to the coefficients of the non-parametric reconstruction of the mass-redshift distribution.
Figure 7: Corner plot showing the posterior distributions of the evolution coefficients for the primary black hole mass ($m_{1}$) using the finer mass binning scheme ($\rm{\Delta M=5M_{\odot}}$). The parameters again correspond to the coefficients $\{p_{0}(m_{k}),p_{1}(m_{k}),p_{2}(m_{k})\}$. The increased number of bins leads to a higher-dimensional parameter space with 42 parameters in total.
Figure 8: Corner plot showing the posterior distributions of the evolution coefficients for the secondary black hole mass ($m_{2}$) using the coarse binning scheme ($\rm{\Delta M=10M_{\odot}}$). The parameters correspond to the Taylor expansion coefficients $\{p_{0}(m_{k}),p_{1}(m_{k}),p_{2}(m_{k})\}$ describing the redshift evolution of the secondary mass distribution. As in the primary mass analysis, the diagonal panels show marginalized one-dimensional posteriors while the off-diagonal panels show joint posterior distributions and parameter correlations.
Figure 9: Corner plot showing the posterior distributions of the evolution coefficients for the secondary black hole mass ($m_{2}$) using the finer mass binning scheme ($\rm{\Delta M=5M_{\odot}}$). The larger number of mass bins leads to a higher-dimensional parameter space with 42 parameters. The corner plot illustrates the posterior constraints and correlations among all inferred coefficients in the hierarchical population model.
@@ Appendix A Posterior Distributions of Evolution Coefficients
The gravitational wave catalogs used in this study are publicly available on Zenodo as part of the LVK data releases: LVK Scientific Collaboration (2021a), LVK Scientific Collaboration (2021b), and LVK Scientific Collaboration (2025).